ING Banner
Home > Astronomy > William Herschel Telescope > Quality Control > WHT optics log

WHT optics log

This page logs numerical results from past Shack-Hartmann tests at the WHT.

They were obtained by analysing (with Dave King's newhart program) the shifts in the x,y positions of the ~ 400 spots between star and lamp (artificial star) exposures. Star and lamp exposures look something like this:

Each data line below gives:

  1. Date yymmdd (of start of night) on which Shack-Hartmann test carried out
  2. Focal station (aux = Cassegrain auxiliary port)
  3. Telescope elevation (deg)
  4. Raw image rms (arcsec)
  5. Raw image 80% diameter (arcsec)
  6. Rms with low-order aberrations subtracted (arcsec)
  7. 80% diameter with low-order aberrations subtracted (arcsec)
  8. Spherical aberration rms (arcsec)
  9. Coma rms (arcsec)
  10. Coma extent (arcsec)
  11. Astigmatism rms (arcsec)
  12. Coma PA (deg) - direction of coma tail measured from 'up' through 'left' on the CCD

For estimates of FWHM, multiply the quoted rms values by 2.35 (assumes the PSF approximates a gaussian).

Orientation: the tests are usually (but not always) carried out with the Shack-Hartmann camera mounted on the WHT folded-Cassegrain ACAM port, with the TEK2 CCD as detector, with instrument = 'OWN' at the TCS and with mount PA = 0. On the CCD, azimuth and elevation directions then lie at approximately 45 deg to the rows and columns, with azimuth increasing to top right and elevation increasing to bottom right on the CCD. A result 'coma PA = 135' then implies that the coma tail points in the direction of decreasing azmiuth.

Increasing mount PA rotates the azimuth and elevation axes anti-clockwise on the TEK2 CCD.

Lines below starting 'CALIBRATE' report the azimuth, elevation and horizontal-collimation errors from a 7-star pointing calibration (CALIBRATE) carried out on the same night.

Changes in telescope pointing, i.e. in the results from the CALIBRATE, are a sensitive diagnostic of misalignment. A coma extent of 1 arcsec will result from any of the following:

  • Translation of M1 by 1.8 mm (pointing error 35 arcsec)
  • Translation of M2 by 1.8 mm (pointing error 27 arcsec)
  • Tilt of M1 by 37 arcsec (pointing error 74 arcsec)
  • Tilt of M2 by 155 arcsec (pointing error 35 arcsec)
For small amounts of translation and tilt, the resulting coma extent and pointing error will scale approximately linearly with translation and tilt.

Coma extent = length of 'teardrop' PSF in Zemax spot diagram.

(Predictions originally provided by Sue Worswick in 1996. Updated by Emilie Lhome and Neil O'Mahony, using Zemax, in July 2017.)

----------------------------------------------------------------------------

9108   Secondary reglued

----------------------------------------------------------------------------

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig
               rms  80%  80%  rms   rms  rms  ext  rms

941117 Aux 88  .11  .38  .28  .08   .05  .05       .05
           58  .10  .36  .27  .07   .02  .05       .04
           30  .10  .37  .26  .07   .02  .06       .04
           83  .10  .37  .26  .07   .02  .04       .06
       (very similar to 12/93 results)

----------------------------------------------------------------------------

9508   proprtional controller for mirror support

----------------------------------------------------------------------------

951007 primary aluminised

951008 wrong spacer, no useful data

----------------------------------------------------------------------------

951120 secondary aluminised

----------------------------------------------------------------------------

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig
               rms  80%  80%  rms   rms  rms  ext  rms

960318 Aux 88  .14  .51  .36  .10   .04  .05  .24  .07
           57  .12  .47  .29  .08   .03  .07  .33  .06
           30  .11  .41  .29  .08   .02  .06  .29  .04
           83  .12  .43  .29  .08   .02  .07  .34  .05

----------------------------------------------------------------------------

961002 primary aluminised, subsequent o/focus images show coma extent
       of approx 1"
961017 mirror support definers tweaked to reduce coma extent < 0.3 arcsec,
       as judged from o/focus images

961122 Aux 82  .12  .43  .32  .09   .01  .05  .26  .06
           52  .12  .43  .32  .09   .04  .06  .30  .03
           82  .12  .44  .32  .09   .04  .05  .25  .05
       Cloudy, test terminated by bad weather.
       Increase of tel AZ changes x,y of spot by -5.5, -3.5 /arcsec
                       EL                        -3.5   5.5  
----------------------------------------------------------------------------

971113 primary aluminised, coma ~ 1 arcsec evident from o/focus images

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig
               rms  80%  80%  rms   rms  rms  ext  rms

971113 Aux 86  .23  .85  .42  .13   .00  .18  .95  .07
           50  .20  .70  .38  .11   .01  .16  .83  .03
           29  .23  .79  .43  .13   .01  .18  .95  .02

971119 primary-mirror axial definers tweaked

971119 Aux 83  .11  .38  .29  .08   .03  .03  .19  .06

(following tests to check optical quality at PF not compromised
by 97119 tweaking)

----------------------------------------------------------------------------

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig
               rms  80%  80%  rms   rms  rms  ext  rms

980414 PF  85  .13  .47  .40  .11   .02  .01  .06  .08  
           74  .26  .74  .65  .22   .02  .02  .13  .13
           64  .16  .57  .43  .12   .02  .04  .18  .10
           50  .14  .50  .48  .13   .00  .04  .22  .03
           down to elev 17 (seeing too poor to use data),
           now coming back up...
           31  .15  .39  .34  .14   .01  .02  .10  .06
           47  .14  .46  .46  .13   .01  .03  .21  .04
           60  .13  .46  .33  .10   .04  .05  .23  .05
           81  .24  .78  .64  .20   .02  .02  .13  .12
           88  .22  .82  .62  .17   .01  .08  .40  .11

----------------------------------------------------------------------------

9809   Primary aluminised (gauges measuring axial position
       of mirror replaced since 1997 aluminising, hence no coma
       expected)

980910 Aux 78                       .06       .31  .07
           62                       .05       .51  .05
           33                       .02       .30  .06
           77                       .01       .38  .12                     

----------------------------------------------------------------------------

990506 1-kg weight fell onto mirror.  Chipped, but no change in CALIBRATE,
       so mirror not moved.

990527 Cass TV images markedly astigamtic, slit has moved

990530 Aux 65  .10  .39  .25  .07   .03  .06  .29  .06

----------------------------------------------------------------------------

000613 Primary aluminised

000616 Aux 77  .11  .37  .28  .09   .01  .05  .26  .05
           50  .15  .36  .32  .12   .06  .06  .29  .02
           31  .14  .40  .34  .11   .06  .07  .37  .03

----------------------------------------------------------------------------

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig
               rms  80%  80%  rms   rms  rms  ext  rms

001108 Aux 75  .14  .47  .34  .10   .01  .07  .46  .06    
           40  .12  .41  .32  .09   .00  .07  .42  .03
(suspected optical problem actually a feature of INGRID)

----------------------------------------------------------------------------

010923 Secondary mount driven to (focus) limit, collision with
       a cable connector introduces large coma.
       Elevation CALIBRATE coefficient has changed by ~ - 40 arcsec.

011029 Secondary remounted.
       Elevation CALIBRATE coefficient back to normal.
       Shack-H mounted at aux port but seeing bad, reduction
       software unable to match spots reliably.

----------------------------------------------------------------------------

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig
               rms  80%  80%  rms   rms  rms  ext  rms

020326 Aux 79  .22  .85  .36  .10   .03  .17  .83  .10
(vector plot similar to that of 011129, so this is probably
residual coma from the 010923 event)

020420 Tilting secondary from handset:
         Transducer readings        1      2        3
         Control-cabinet channels   7      8        9
         Correspond to            pivot  Q motor  P motor
         Pivot is in elevation direction.
         Q-, P+ on handset increases transducer readings 2,3
         assumed to indicate movement in elevation.
         Q and P speeds are 4.0 and 2.1 microns/sec. 
         TCS displays of transducers out of range at +-100, 
         so used LCD displays of LVDTs instead. 
         
Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig Coma
               rms  80%  80%  rms   rms  rms  ext  rms   PA
020420 At start of night, star image slightly comatic on ShackH TV view.

020420 Aux 81  .17  .59  .34  .10   .05  .12  .81  .02    -6
           36  .20  .69  .35  .11   .08  .13  .83  .05   -14
Move secondary transducers P and Q 50 microns and back.
           67                       .08  .12  .76  .06   -11
Secondary LVDTs now read -38 85 69 on LCD display.
Tilt secondary in elev to -74 135 119 i.e. 50 microns each on Q and P
           81                       .07  .12  .81  .03    -6
Tilt secondary in elev to -101 185 169 on LCDs i.e. another 50 microns.
           77                       .10  .10  .67  .03    -6
Tilt secondary in elev to -154 285 269 on LCDs i.e. another 100 microns.
           70                       .09  .06  .34  .06    47
Tilt secondary in elev to -199 285 469 on LCDs i.e. another 200 microns
on Q.
           59                       .07  .04  .22  .05   133
           59                       .06  .03  .18  .06   159
           79                       .06  .06  .37  .08  -172
           79                       .06  .07  .38  .10   161
           35                       .07  .10  .59  .13   140
           35                       .08  .05  .25  .09   118
           82  .18  .62  .47  .14   .02  .04  .23  .10   138
NB PA of coma (from sharp to blunt end) on CCD is measured clockwise
from vertical.  Stars move at PA -45 when elevation increased, PA -135 when
azimuth increased.
During reduction, inner and outer spots masked out but vanes invisible 
(up-down, left-right) so not masked.
No evidence after tweaking for distortions of in-focus or out-of-focus
images on ShackH TV.
Net movement was +200 microns on P, +400 on Q, i.e. a tilt ~ 80 arcsec
(+- 20%).
A tilt of this magnitude should cause coma 0.6 arcsec, which is
consistent.  NB we might have simply compensated for translation by
introducing tilt, which would degrade image quality off-axis,
but this will be < 0.2 arcsec at 2' radius (e.g. INGRID).

020420 UES CALIBRATE coeffs AZ EL HC have changed from -21 -140 70
       to -25 -90 130.  A change of 50 arcsec in the EL term
       corresponds to secondary tilt 110 arcsec, broadly consistent.

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
               rms  80%  80%  rms   rms  rms  ext  rms   PA
020425 Aux 66  .09  .32  .28  .08   .01  .05  .24  .02   123    
Checking no change after top-end flipped.

----------------------------------------------------------------------------

021218 Primary mirror replaced after first aluminising for 2.5 years

021218 Aux 79  .15  .49  .38  .13   .00  .06  .31  .05    79
       Aux 45  .14  .47  .41  .13   .00  .05  .25  .02   135 

----------------------------------------------------------------------------

030409 o/focus image pair taken at aux port by AZ, no aberrations

----------------------------------------------------------------------------

0411** secondary aluminised

041123 pointing ~ 1' out in elevation, focus at LIRIS, aux and GRACE
       prob increased by 0.3 mm
       o/focus image taken on LIRIS, aberrations < 0.5 arcsec
       Shack-Hartmann OK:

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
               rms  80%  80%  rms   rms  rms  ext  rms   PA
041123 Aux 84  .12  .42  .28  .09   .05  .03  .18  .06   137
        
       The support of the secondary mirror was apparently too rigid, 
       due to a misplaced washer on one of the springs.  This could
       affect pointing due to flexure of the mirror backplane being
       picked up by the LVDT transducers and fed back to the TCS,
       without affecting optical alignment.
       Support adjusted.

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
               rms  80%  80%  rms   rms  rms  ext  rms   PA
041124 Aux 86  .12  .41  .35  .10   .05  .02  .13  .06    73
041124 Aux 57  .14  .46  .35  .12   .03  .05  .27  .04   116

041202 Primary aluminised

041211 Aux 76  .08  .28  .26  .07   .03  .01  .09  .02  -159

----------------------------------------------------------------------------

070131 Primary aluminised

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
               rms  80%  80%  rms   rms  rms  ext  rms   PA
070202 Aux 75  .13  .48  .36  .10   .05  .04  .21  .05   -88
           70  .15  .53  .38  .11   .03  .07  .36  .07   -74
           46  .14  .52  .47  .13   .04  .05  .24  .03  -137
           33  .10  .38  .25  .08   .03  .03  .18  .05   174

----------------------------------------------------------------------------

070405 Test of optical quality at the previously-unused broken-Cass focus, 
       in anticipation of the possible deployment there of the HARPS-N
       fibre feed.
       
       Out-of-focus white-light images were obtained (the Shack-Hartmann 
       camera could not be mounted) at telescope focus 92.0 and 104.0 mm (best
       focus +- 5 mm), elevation 70 deg.  Results:
         COMA:            < 0.3 arcsec (FWHM)
         SPH. ABERRATION: < 0.2 arcsec (diameter of disc of least confusion)
         ASTIGMATISM:     < 0.3 arcsec (diameter at best mean focus)
       These limits are consistent with the values measured using the 
       Shack-Hartmann camera at the Cass auxiliary port 2 months earlier
       (070202 above), each aberration ~ 0.1 arcsec FWHM.
       
       On an image of the globular cluster M3, the images are round,
       with FWHM 1.3 arcsec, and no apparent changes in form across
       the 60-arcsec field of the camera (0.06 arcsec/pixel).

       An image taken at elevation 20 deg showed the (expected) ~ 2-arcsec
       dispersion in the elevation direction, but otherwise nothing 
       unusual.

----------------------------------------------------------------------------

091103 Primary aluminised (no subsequent Shack-Hartmann, because of bad weather)

----------------------------------------------------------------------------

120208 Primary aluminised

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
               rms  80%  80%  rms   rms  rms  ext  rms   PA

120209 Aux 82  .13  .43  .33  .09   .03  .01  .04  .08  -151
           34  .12  .45  .34  .09   .01  .02  .10  .08   156

----------------------------------------------------------------------------

140519 PF  87  .16  .55  .35  .10   .03  .10  .51  .07   -43
           78  .16  .56  .25  .08   .07  .09  .46  .08   -44
           63  .14  .50  .38  .10   .05  .08  .35  .01   -54
           51  .15  .55  .36  .09   .02  .11  .51  .02   -63
           47  .12  .43  .28  .08   .01  .09  .43  .04   -56
           25  .16  .57  .29  .08   .05  .11  .62  .05   -27

----------------------------------------------------------------------------

140915 Shack-Hartmann tests at bare prime focus

----------------------------------------------------------------------------

160519 Primary aluminised.  Transverse-definer pad T3 (on side of M1)
came unglued, re-glued.

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
yymmdd         rms  80%  80%  rms   rms  rms  ext  rms   PA

160520 Aux 71  .28  .91  .30  .09   .02  .25 1.38  .09  -138 (mt PA = 0)

       => coma in azimuth

       CALIBRATE AZ, EL, HC = -22, 6, 143 
       (cf median ~ -25, -22, 100 Jan 2015 - mid-May 2016)

160521 Aux 49  .26  .92  .35  .11   .08  .21 1.16  .06    40 (mt PA = -180)
           33  .29 1.04  .34  .11   .05  .25 1.44  .10    37 ( " " )

       => coma in azimuth (+- few deg)

       CALIBRATE AZ, EL, HC = -13, -46, 149

160523 Aux 74  .35 1.18  .36  .10   .02  .32 1.70  .08  -130 (mt PA = -3)

160524 M1 transverse definer T3 adjusted

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
yymmdd         rms  80%  80%  rms   rms  rms  ext  rms   PA

       Aux 59  .31 1.12  .34  .09   .00  .28 1.53  .09  -136 (mt PA = 0)
           52  .28  .99  .37  .11   .03  .25 1.27  .08  -138 ( " " )

       => coma still in azimuth, slightly worse

       CALIBRATE AZ, EL, HC = -28, -21, 165

160526 M1 translated 2.8 mm in GRACE-GHRIL direction

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
yymmdd         rms  80%  80%  rms   rms  rms  ext  rms   PA

       Aux 74  .27  .93  .56  .16   .10  .07  .45  .19  -178 (mt PA = 0)
       
           38  .17  .57  .44  .13   .06  .09  .52  .03   168 ( " " )
       Later same night:
           77  .28 1.07  .69  .20   .06  .06  .36  .17  -168 ( " " )

       => coma much reduced, but small astigmatism at high elevation

       CALIBRATE AZ, EL, HC = -25, -26, 115 
           
----------------------------------------------------------------------------

160613 Further investigations of aberrations at Cass
       All data taken with mount PA = 0
       
       CALIBRATE AZ, EL, HC = -19, -31, 110

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
yymmdd         rms  80%  80%  rms   rms  rms  ext  rms   PA

160613 Aux 84  .10  .37  .29  .08   .01  .04  .25  .04   166
           35  .13  .44  .33  .09   .04  .06  .36  .04   170
      
       I.e. image quality now very good, although M1 has not been
       touched since 160526 (although it might have been relaxing to 
       its current position under repeated telescope movements).

       Point to zenith.  Definer T3 relaxed.  

       Aux 71  .16  .57  .41  .12   .05  .05  .35  .08   180
           36  .16  .52  .33  .10   .07  .05  .31  .08   179

       Point to zenith.  Definer T3 restored to start-of-night state.
  
       CALIBRATE AZ, EL, HC = -32, -25, 115

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
yymmdd         rms  80%  80%  rms   rms  rms  ext  rms   PA

160613 Aux 76  .13  .47  .32  .09   .03  .06  .35  .07   178
           51  .16  .56  .45  .11   .03  .04  .26  .09   160
           28  .10  .34  .32  .10   .03  .01  .07  .02    76
           51  .13  .46  .32  .09   .01  .04  .21  .08  -159
           77  .12  .45  .36  .10   .00  .04  .37  .06  -159      

Measurements on next 2 nights (no adjustments to M1):
      
160614 two CALIBRATEs during night:

       CALIBRATE AZ, EL, HC = -18, -33, 108

       CALIBRATE AZ, EL, HC = -25, -34, 111

160614 Aux 72  .13  .47  .30  .08   .01  .06  .37  .07  -154  
           20  .12  .46  .34  .10   .00  .07  .42  .04   172
           30  .10  .34  .32  .09   .01  .04  .18  .02   118
           48  .18  .63  .40  .12   .04  .05  .26  .12   110

160615 Aux 51  .10  .34  .30  .08   .01  .00  .02  .05   110
           72  .14  .47  .40  .12   .03  .06  .30  .04   141
           34  .10  .37  .29  .08   .02  .04  .27  .04   180
           48  .10  .34  .32  .09   .01  .02  .11  .02   -84

Re-ordering the results of the above 13 tests by elevation:
      
Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
yymmdd         rms  80%  80%  rms   rms  rms  ext  rms   PA

           20  .12  .46  .34  .10   .00  .07  .42  .04   172
           28  .10  .34  .32  .10   .03  .01  .07  .02    76
           30  .10  .34  .32  .09   .01  .04  .18  .02   118
           34  .10  .37  .29  .08   .02  .04  .27  .04   180

           48  .10  .34  .32  .09   .01  .02  .11  .02   -84
           48  .18  .63  .40  .12   .04  .05  .26  .12   110
           51  .10  .34  .30  .08   .01  .00  .02  .05   110
           51  .13  .46  .32  .09   .01  .04  .21  .08  -159
           51  .16  .56  .45  .11   .03  .04  .26  .09   160

           72  .13  .47  .30  .08   .01  .06  .37  .07  -154  
           72  .14  .47  .40  .12   .03  .06  .30  .04   141
           76  .13  .47  .32  .09   .03  .06  .35  .07   178
           77  .12  .45  .36  .10   .00  .04  .37  .06  -159      
 
Conclusions from measurements over the last 3 nights:

   (1) At a given elevation, the derived values for each of
       spherical-aberration, coma and astigmatism rms are consistent 
       from night to night within ~ +-0.03 arcsec.

   (2) Astigmatism is consistently lower at elevation ~ 30 deg than
       at higher elevations.  Coma may be slightly higher near the zenith than
       at lower elevations.

   (3) The high-elevation astigmatism has decreased from the night 
       160613 to the night 160615.  The significance of this is not
       high, but the change might reflect relaxation of M1 after the tweaking
       of definer T3 on 160613.

   (4) The quality of the delivered PSF is currently similar to that
       delivered over most of the past 20 years, albeit with coma and 
       astigmatism towards the upper end of the usual range:
                            
                            Typical,         June 2016
                            1990s - 2012     (high elevation)
                            (arcsec)         (arcsec)

       Raw rms:             .12 - .14        .13

       Sph aberration rms:  .01 - .04        .02
       Coma rms:            .04 - .07        .06
       Coma extent:         .10 - .30        .35
       Astigmatism:         .03 - .07        .06

       The delivered FWHM is 2.35 * raw rms ~ 0.33 arcsec.

----------------------------------------------------------------------------

160718 Prior to removal of the WHT top-end flip ring on 160719

       All data for 160718 taken with mount PA = 0

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
yymmdd         rms  80%  80%  rms   rms  rms  ext  rms   PA

160718 Aux 80  .15  .52  .42  .11   .01  .04  .27  .08   152
           48  .11  .40  .29  .08   .00  .06  .36  .05   176
           37  .10  .38  .31  .08   .01  .04  .23  .05  -164
           23  .11  .41  .28  .08   .01  .07  .44  .03   177           

       CALIBRATE AZ, EL, HC = -24, -29, 112

160719 WHT top-end flip ring removed, and replaced the next day (test   
       for WEAVE).

160721 Aux 80  .11  .38  .34  .09   .00  .03  .14  .05   -67 (mt PA = -125)
                                                      => 168 (mt PA = 0) 

           59  .11  .41  .32  .09   .01  .03  .15  .06  -106 (mt PA = -84)
                                                      => 170 (mt PA = 0)

           31  .11  .41  .30  .08   .01  .06  .34  .05  -147 (mt PA = -58)
                                                      => 157 (mt PA = 0)

           27  .10  .35  .29  .08   .02  .04  .25  .04   178 (mt PA = -7)
                                                      => 171 (mt PA = 0)

       CALIBRATE AZ, EL, HC = -34, -38, 110

Conclusion: No degradation in optical quality after removal and 
replacement of flip ring.  Sizes of coma, astigmatism and spherical 
aberration fall well within ranges observed over the last two decades.

----------------------------------------------------------------------------
Test at prime focus (with PF corrector in light path)

Date   Foc El  Raw  Raw  Sub  Sub   Sph    Coma   Astig  Coma
yymmdd         rms  80%  80%  rms   rms  rms  ext  rms   PA

160921 PF  60  .11  .39  .33  .09   .00  .05  .24  .04   -109 

----------------------------------------------------------------------------

170407 After removal and replacement of the WHT top-end ring, both 
       intra-/extra-focal imaging with ACAM, and comparison of before and 
       after CALIBRATE index errors indicate coma extent < 0.3 arcsec.
       (A Shack-Hartmann test was not possible because the ACAM port was
       needed for science.)

----------------------------------------------------------------------------

170619 After removal and replacement of the WHT top-end ring, ACAM images
       taken in 0.6 arcsec seeing have excellent circular symmetry, no sign 
       of any coma.  The alignment of pupil and central hole on an out-of-focus
       image of a star confirms coma extent < 0.6 arcsec, and comparison of 
       before and after CALIBRATE index errors indicates coma extent 
       < 0.4 arcsec.
       (The Shack-Hartmann camera was not available.)

----------------------------------------------------------------------------

Back to WHT optics page.



Top | Back

Contact:  (WHT Manager)
Last modified: 03 July 2017